# Central Dogma Transformer II An AI Microscope for Understanding Cellular Regulatory Mechanisms

### v1

@@ Abstract
Current biological AI models lack interpretability—their internal representations do not correspond to biological relationships that researchers can examine. Here we present CDT-II, an “AI microscope” whose attention maps are directly interpretable as regulatory structure. By mirroring the central dogma in its architecture, each attention mechanism corresponds to a specific biological relationship: DNA self-attention for genomic relationships, RNA self-attention for gene co-regulation, and DNA-to-RNA cross-attention for transcriptional control. Using only genomic embeddings and raw per-cell expression, CDT-II enables experimental biologists to observe regulatory networks in their own data. Applied to K562 CRISPRi data, CDT-II predicts perturbation effects (per-gene mean $r=0.84$) and recovers the GFI1B regulatory network without supervision (6.6-fold enrichment, $P=3.5\times 10^{-17}$). Two distinct attention mechanisms converge on an RNA processing module ($P=1\times 10^{-16}$). CDT-II establishes mechanism-oriented AI as an alternative to task-oriented approaches, revealing regulatory structure rather than merely optimizing predictions.
@@ Introduction
The central dogma1 describes cellular information flow across three molecular layers: DNA encodes genetic instructions, RNA transmits and regulates this information, and proteins execute cellular functions. Artificial intelligence has transformed our ability to model each layer individually2–7, yet a fundamental limitation remains: current models lack interpretability. Their internal representations do not correspond to biological entities or relationships that researchers can examine and validate. The goal of biological research, however, is not merely to predict cellular responses but to understand the mechanisms that produce them. What biology needs is not task-oriented AI optimizing predictions, but mechanism-oriented AI revealing regulatory structure.
To address this limitation, we reframe biological AI as an “AI microscope” whose attention maps are directly interpretable as regulatory structure. The Central Dogma Transformer8 implements this principle by mirroring the central dogma in its architecture: DNA self-attention captures genomic relationships, RNA self-attention reflects gene co-regulation, and DNA-to-RNA cross-attention models transcriptional control. Because each attention mechanism corresponds to a specific biological relationship, the resulting maps provide direct readouts of regulatory organization. This architectural inductive bias fundamentally changes the meaning of prediction accuracy: unlike models optimizing prediction as an end goal, CDT-II uses prediction accuracy as evidence that the model has learned correct regulatory structure. As training progresses, improving predictions and discovering biology become the same objective.
@@
While CDT v1 required three pre-trained language models, CDT-II reduces the input requirements to genomic sequence embeddings and raw per-cell expression, enabling experimental biologists to observe regulatory networks in their own data.
We validated CDT-II on a large-scale K562 CRISPRi screen9,10, holding out five perturbation targets entirely—including GFI1B, a master transcriptional regulator with hundreds of experimentally characterized trans-regulatory targets10—to test whether regulatory networks could be recovered from unsupervised attention patterns. Here we report four principal findings. First, CDT-II achieves a validation Pearson correlation of 0.64 (per-gene mean $r=0.84$) without requiring RNA or protein language models. Second, attention analysis recovers GFI1B’s trans-regulatory network with 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$), demonstrating unsupervised discovery of a known regulatory program. Third, two distinct attention mechanisms—RNA self-attention and DNA-to-RNA cross-attention—converge on the same RNA processing module ($P=1\times 10^{-16}$), revealing hierarchical regulation across central dogma layers. Fourth, an ablation study establishes that gene set quality, determined by cross-dataset reproducibility, governs model resolution. Together, these results demonstrate that CDT-II’s attention maps provide biologically meaningful observations of regulatory networks.
@@ Results
CDT-II implements a two-modality architecture that follows the directional logic of the central dogma (Fig. 1). Genomic DNA embeddings, generated by Enformer3 from sequences centered on perturbation sites, are represented as a $896\times 3{,}072$ matrix (896 genomic bins across the 196-kilobase input window, each with a 3,072-dimensional embedding). Per-cell RNA expression values for 2,361 genes are encoded by a RawExpressionEncoder that combines learned gene identity embeddings with projected expression values, producing gene-level representations of dimension 512.
Each modality first passes through self-attention layers11 that capture internal structure: two layers for DNA (learning relationships between genomic positions) and one layer for RNA (learning co-regulatory relationships between genes). A cross-attention layer then models transcriptional regulation, with RNA representations serving as queries and DNA representations as keys and values—enabling the model to learn which genomic regions are relevant for each gene’s regulation. All attention operations use 8 heads and preserve the original dimensionality, ensuring that attention maps remain directly interpretable as gene-by-gene or gene-by-position matrices.
After cross-attention, a Virtual Cell Embedding (VCE)8 module integrates DNA and RNA representations into a unified 512-dimensional cell-state vector. This integration uses learned attention pooling, which weights each gene’s contribution based on its relevance. A task layer then projects this VCE to predict log2 fold changes for all 2,361 genes simultaneously. The complete model contains approximately 21 million parameters (see Online Methods for training details).
The dataset comprised 15,657 K562 cells from the CRISPRi screen10: 8,250 cells from TSS perturbations (27 genes) and 7,407 cells from SNP perturbations (420 loci). Of these, 13,620 cells were used for training and 2,037 cells were reserved for validation. Training included 6,213 TSS cells (22 genes) and all 7,407 SNP cells. Validation comprised five genes held out entirely at the gene level: GFI1B, CD52, TFRC, CD44, and TNFSF9. This gene-level holdout ensures that validation performance reflects generalization to entirely unseen perturbation targets—a stringent test of the model’s ability to learn transferable regulatory patterns rather than memorize gene-specific effects.
@@
Figure 1: CDT-II architecture and interpretable attention maps. Unlike conventional deep learning models whose internal representations are opaque, CDT-II produces attention maps that directly correspond to biological relationships. CDT-II processes two input modalities following the central dogma. Left branch: Genomic DNA sequences centered on perturbation loci ($\pm$98 kb) are encoded by Enformer, projected to a common dimension, and processed by two DNA self-attention layers that capture genomic relationships between positions. Right branch: Per-cell RNA expression values from scRNA-seq data are encoded by the RawExpression Encoder, projected, and processed by one RNA self-attention layer that captures gene co-regulatory relationships. Center: A DNA-to-RNA cross-attention layer models transcriptional control, with RNA representations as queries and DNA representations as keys and values. The Virtual Cell Embedder integrates both modalities through attention pooling into a unified cell-state vector, which the Task Head projects to predict perturbation effects for all 2,361 genes. Right panels: Each attention mechanism produces an interpretable map—DNA self-attention reveals genomic relationships (blue), RNA self-attention reveals gene co-regulation networks (orange), and cross-attention reveals transcriptional control patterns (purple). These maps constitute the primary output of the “AI microscope,” enabling direct observation of regulatory structure.
@@
Our initial attempt used 9,335 genes detected in at least 10% of cells in the CRISPRi dataset10, applying only a single-dataset expression threshold. This configuration ($\sim$54 million parameters) plateaued at a validation Pearson $r$ of 0.36, and attention maps exhibited collapse—a diffuse, uninformative pattern indicating that the model could not resolve regulatory structure from the noisy gene set.
Based on this result, we hypothesized that genes lacking reproducibility across independent experiments introduced noise that degraded model performance. We therefore applied a cross-dataset filter, retaining only the 2,360 genes detected as expressed in two independent K562 CRISPRi screens10,12, and added GFI1B as a critical test case (2,361 total, $\sim$21 million parameters). The effect was immediate: with reduced noise, the learning curve improved steadily throughout training, reaching $r=0.64$ with training r=0.65r=0.65—indicating minimal overfitting.
The contrast is stark: nearly quadrupling the gene count degraded performance by 43%, while filtering to cross-dataset reproducible genes yielded a 78% improvement. This establishes a counterintuitive but practically important principle: gene set quality, not quantity, determines the resolution of the AI microscope. Cross-dataset reproducibility filtering removes genes whose perturbation responses are dominated by technical or biological noise, enabling the model to learn sharper regulatory relationships.
@@
To evaluate how well CDT-II captures the trans-regulatory “fingerprint” of each perturbation, we averaged predictions and observations across cells for each held-out gene and computed Pearson $r$ on these mean trans-effect profiles: GFI1B ($r=0.86$), TNFSF9 ($r=0.86$), TFRC ($r=0.83$), CD44 ($r=0.81$), and CD52 ($r=0.64$), yielding a mean $r$ of 0.84. This high pseudo-bulk correlation demonstrates that CDT-II faithfully captures the transcriptome-wide pattern of perturbation effects—the regulatory “fingerprint” that distinguishes one perturbation from another.
To contextualize these correlations, we examined the empirical reproducibility of CRISPRi perturbation effects across independent studies. Morris et al. found that only 56% of GFI1B trans-targets identified in their STING-seq screen replicated in the independent Gasperini dataset10,12, and direct comparison of pseudo-bulk effect sizes between the two K562 CRISPRi studies yields $r\approx 0.79$. CDT-II’s pseudo-bulk correlation ($r=0.84$) approaches this empirical ceiling, indicating that the model captures essentially all reproducible biological signal. The lower cell-level correlation ($r=0.64$) reflects stochastic variation inherent to single-cell measurements—noise that is averaged out at the pseudo-bulk level but represents a fundamental limit of single-cell prediction.
Trans-target prediction provided a stringent test of CDT-II’s ability to capture regulatory wiring. Among the top 50 predicted trans-targets, GFI1B showed substantial overlap with experimentally determined top targets (versus 2.1% expected by chance), while CD52 showed similar enrichment. Notable GFI1B trans-targets correctly predicted include G0S2, TXNIP, and JUN—genes involved in cell cycle arrest, oxidative stress response, and transcriptional regulation, respectively, consistent with GFI1B’s known functions in hematopoietic differentiation.
@@
Figure 3: CDT-II predicts cell-level perturbation effects. (A) Overall scatter plot of predicted versus actual log2 fold changes across all validation cells (n=2,037n=2{,}037 cells $\times$ 2,361 genes), showing Pearson $r=0.64$. Each point represents a cell-gene pair, with density indicated by color intensity. (B) Per-gene prediction performance for five validation genes (GFI1B, TNFSF9, TFRC, CD44, CD52), with pseudo-bulk correlations (mean across cells) ranging from 0.64 to 0.86 (mean $r=0.84$). (C) Trans-effect profile for GFI1B: mean predicted versus mean actual effects across all 2,361 genes, showing strong recovery of the perturbation-specific regulatory signature ($r=0.86$).
@@
The attention analysis follows the central dogma flow. DNA self-attention (Fig. 4A) reveals how the model learns relationships between genomic positions within the 196 kb window centered on the perturbation site. DNA-to-RNA cross-attention (Fig. 4B) shows how each gene attends to specific genomic regions, capturing transcriptional regulatory relationships. RNA self-attention (Fig. 4C–D) captures co-regulatory relationships between genes, revealing the gene regulatory network structure.
The RNA self-attention matrix $2{,}361\times 2{,}361$] captures learned co-regulatory relationships between genes. Extracting the GFI1B query row—representing which genes GFI1B “attends to” during inference—revealed a biologically coherent ranking. Cell cycle regulators including CDCA8, CDC20, KIF2C, and KIF14 appeared in the top ranks, consistent with GFI1B’s established role in cell cycle control during hematopoietic differentiation13.
Visualizing the top 5% of attention edges as a network (Fig. 5) reveals GFI1B as a hub gene with multiple connection types: downstream targets (genes that GFI1B regulates), upstream regulators (genes that influence GFI1B), and bidirectional relationships suggesting reciprocal regulatory dynamics. This bidirectional structure captures the feedback loops characteristic of transcription factor networks.
To quantify the correspondence between attention-derived and experimentally determined regulatory networks, we ranked all 2,361 genes by GFI1B attention weight and, independently, by absolute experimental effect size (mean $|\log_{2}\text{FC}|$ upon GFI1B perturbation10), then measured overlap between the top $N$ genes from each ranking. At top 100, 28 genes appeared in both lists, representing a 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$ by hypergeometric test). This demonstrates that CDT-II’s attention maps recover known regulatory relationships without supervision.
@@
Figure 4: Attention maps reveal regulatory structure. (A) DNA self-attention patterns across the 196 kb genomic window for two layers. (B) DNA-to-RNA cross-attention profiles for five TSS perturbations.
@@
Figure 4 (continued). (C) RNA self-attention per-head analysis for GFI1B, showing that different attention heads capture distinct aspects of the regulatory network. (D) RNA self-attention 2D heatmap showing gene-gene co-regulatory relationships centered on GFI1B and its top-attended genes. (E) Cell-to-cell attention variation. Pairwise correlations of GFI1B attention vectors across validation cells show mean $r=0.75$, indicating that while core network structure is preserved, attention weights are modulated by each cell’s expression state.
@@
We applied Louvain community detection14 to networks derived from each mechanism separately. From the RNA self-attention matrix, we identified community C1 (679 genes) with highly significant enrichment for RNA binding (Gene Ontology15 molecular function; $P=5\times 10^{-16}$) and metabolism of RNA (Reactome16; $P=2\times 10^{-9}$). From the DNA-to-RNA cross-attention matrix—where gene similarity was computed based on shared attention to the same genomic regions—we identified community C3 (165 genes) with even stronger enrichment for RNA binding ($P=1\times 10^{-16}$), spliceosome (KEGG17; $P=4\times 10^{-10}$), and mRNA splicing (Reactome16; $P=2\times 10^{-9}$).
The convergence between these independently derived modules was striking: 132 of 165 cross-attention genes (80%) were contained within the 679-gene self-attention community. This core module of 132 genes spans multiple levels of RNA processing: spliceosome components (SF3A3, SF3B4, PRPF3, SNRNP27, SNRPG, BCAS2, RBM8A), RNA modification and decay factors (YTHDF2 for m6A reading, ADAR for RNA editing, EXOSC10 for exosome-mediated degradation, TUT4 for uridylation, XPO1 for nuclear export), chromatin regulators (HDAC1, KDM1A/LSD1, ARID1A of the SWI/SNF complex), transcription machinery (CDC73 of the PAF complex, TAF12 of TFIID, TRIM33), and translation initiation factors (EIF2B3, EIF2B4, EIF4G3).
@@
That both mechanisms independently identify the same RNA processing module indicates that these genes are coordinately regulated at two distinct levels of the central dogma: they share transcriptional regulatory inputs (detected by cross-attention) and participate in the same post-transcriptional feedback networks (detected by self-attention) (Fig. 6). Notably, while the architecture provides separate attention mechanisms, this convergence was not explicitly encouraged—it emerged from training on perturbation data alone.
Figure 5: GFI1B regulatory subnetwork derived from RNA self-attention. Network visualization of GFI1B regulatory relationships derived from the top 5% of attention edges. GFI1B (red, center) functions as a hub gene connected to downstream targets (blue), upstream regulators (orange), and genes with bidirectional relationships (purple) suggesting reciprocal regulatory dynamics. Edge thickness is proportional to attention weight. The network structure reveals both the genes that GFI1B regulates and the feedback loops characteristic of transcription factor networks. Bottom text shows network statistics: out-degree (downstream targets) and in-degree (upstream regulators) from the full attention matrix.
@@
Figure 6: Convergent attention reveals an RNA processing module. (A) GO enrichment comparison showing both attention mechanisms identify “RNA Binding” as the top enriched term. RNA self-attention community C1 (679 genes) and DNA-to-RNA cross-attention community C3 (165 genes) show highly significant enrichment for RNA processing functions. (B) Venn diagram showing 132-gene core overlap (80% of cross-attention community). Core genes span spliceosome components, RNA processing factors, chromatin regulators, and transcription machinery.

### v2

@@ Abstract
[…]
@@ Introduction
The central dogma1 describes cellular information flow across three molecular layers: DNA encodes genetic instructions, RNA transmits and regulates this information, and proteins execute cellular functions. Artificial intelligence has transformed our ability to model each layer individually2–7, yet a fundamental limitation remains: current models lack interpretability8,9. Their internal representations do not correspond to biological entities or relationships that researchers can examine and validate. The goal of biological research, however, is not merely to predict cellular responses but to understand the mechanisms that produce them. What biology needs is not task-oriented AI optimizing predictions, but mechanism-oriented AI revealing regulatory structure.
To address this limitation, we reframe biological AI as an “AI microscope” whose attention maps are directly interpretable as regulatory structure. The Central Dogma Transformer10 implements this principle by mirroring the central dogma in its architecture: DNA self-attention captures genomic relationships, RNA self-attention reflects gene co-regulation, and DNA-to-RNA cross-attention models transcriptional control. Because each attention mechanism corresponds to a specific biological relationship, the resulting maps provide direct readouts of regulatory organization. This architectural inductive bias fundamentally changes the meaning of prediction accuracy: unlike models optimizing prediction as an end goal, CDT-II uses prediction accuracy as evidence that the model has learned correct regulatory structure. As training progresses, improving predictions and discovering biology become the same objective.
@@
While CDT v1 required three pre-trained language models, CDT-II reduces the input requirements to genomic sequence embeddings and raw per-cell expression, enabling experimental biologists to observe regulatory networks in their own data. CDT-II is designed as an integration platform rather than a standalone model: pre-trained language models provide the best available representations of each molecular layer, while task-driven training on experimental data organizes this knowledge into interpretable regulatory maps. The DNA embedding component is deliberately modular, allowing it to be upgraded as genomic foundation models improve. Because each attention mechanism corresponds to a specific biological relationship, CDT-II’s outputs generate experimentally testable hypotheses—creating a feedback cycle in which experimental validation improves the model, and the improved model generates sharper hypotheses for the next round of experiments.
We validated CDT-II on a large-scale K562 CRISPRi screen11,12, holding out five perturbation targets entirely—including GFI1B, a master transcriptional regulator with hundreds of experimentally characterized trans-regulatory targets12—to test whether regulatory networks could be recovered from unsupervised attention patterns. Here we report five principal findings. First, CDT-II achieves a validation Pearson correlation of 0.64 (per-gene mean $r=0.84$) without requiring RNA or protein language models. Second, attention analysis recovers GFI1B’s trans-regulatory network with 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$), demonstrating unsupervised discovery of a known regulatory program. Third, two distinct attention mechanisms—RNA self-attention and DNA-to-RNA cross-attention—independently identify an overlapping RNA processing module (80% gene overlap) with highly significant functional enrichment (RNA binding, $P=1\times 10^{-16}$), revealing hierarchical regulation across central dogma layers. Fourth, an ablation study establishes that gene set quality, determined by cross-dataset reproducibility, governs model resolution. Fifth, systematic comparison against ENCODE K562 regulatory annotations reveals that cross-attention autonomously focuses on known regulatory elements, including DNase hypersensitive sites and CTCF binding sites, across all five held-out genes. Together, these results demonstrate that CDT-II’s attention maps provide biologically meaningful observations of regulatory networks.
@@ Results
CDT-II implements a two-modality architecture that follows the directional logic of the central dogma (Fig. 1). Genomic DNA embeddings, generated by Enformer3 from sequences centered on perturbation sites, are represented as a $896\times 3{,}072$ matrix (896 genomic bins spanning approximately 115 kilobases at 128-base-pair resolution, each with a 3,072-dimensional embedding). Per-cell RNA expression values for 2,361 genes are encoded by a RawExpressionEncoder that combines learned gene identity embeddings with projected expression values, producing gene-level representations of dimension 512.
Each modality first passes through self-attention layers13 that capture internal structure: two layers for DNA (learning relationships between genomic positions) and one layer for RNA (learning co-regulatory relationships between genes). A cross-attention layer then models transcriptional regulation, with RNA representations serving as queries and DNA representations as keys and values—enabling the model to learn which genomic regions are relevant for each gene’s regulation. All attention operations use 8 heads and preserve the original dimensionality, ensuring that attention maps remain directly interpretable as gene-by-gene or gene-by-position matrices.
After cross-attention, a Virtual Cell Embedding (VCE)10 module integrates DNA and RNA representations into a unified 512-dimensional cell-state vector. This integration uses learned attention pooling, which weights each modality’s elements based on their relevance. A two-layer task head then projects this VCE to predict log2 fold changes for all 2,361 genes simultaneously. The complete model contains approximately 21 million parameters (see Online Methods for training details).
The dataset comprised 15,657 K562 cells from the CRISPRi screen12: 8,250 cells from TSS perturbations (27 genes) and 7,407 cells from SNP perturbations (420 loci). Of these, 13,620 cells were used for training and 2,037 cells were reserved for validation. Training included 6,213 TSS cells (22 genes) and all 7,407 SNP cells. Validation comprised five genes held out entirely at the gene level: GFI1B, CD52, TFRC, CD44, and TNFSF9. This gene-level holdout ensures that validation performance reflects generalization to entirely unseen perturbation targets—a stringent test of the model’s ability to learn transferable regulatory patterns rather than memorize gene-specific effects.
@@
Figure 1: CDT-II architecture and interpretable attention maps. Unlike conventional deep learning models whose internal representations are opaque, CDT-II produces attention maps that directly correspond to biological relationships. CDT-II processes two input modalities following the central dogma. Left branch: Genomic DNA sequences centered on perturbation loci ($\pm$57 kb) are encoded by Enformer, projected to a common dimension, and processed by two DNA self-attention layers that capture genomic relationships between positions. Right branch: Per-cell RNA expression values from scRNA-seq data are encoded by the RawExpressionEncoder, projected, and processed by one RNA self-attention layer that captures gene co-regulatory relationships. Center: A DNA-to-RNA cross-attention layer models transcriptional control, with RNA representations as queries and DNA representations as keys and values. The Virtual Cell Embedder integrates both modalities through attention pooling into a unified cell-state vector, which the Task Head projects to predict perturbation effects for all 2,361 genes. Right panels: Each attention mechanism produces an interpretable map—DNA self-attention reveals genomic relationships (blue), RNA self-attention reveals gene co-regulation networks (orange), and cross-attention reveals transcriptional control patterns (purple). These maps constitute the primary output of the “AI microscope,” enabling direct observation of regulatory structure.
@@
Our initial attempt used 9,335 genes detected in at least 10% of cells in the CRISPRi dataset12, applying only a single-dataset expression threshold. This configuration ($\sim$54 million parameters) plateaued at a validation Pearson $r$ of 0.37, and attention maps exhibited collapse—a diffuse, uninformative pattern indicating that the model could not resolve regulatory structure from the noisy gene set.
Based on this result, we hypothesized that genes lacking reproducibility across independent experiments introduced noise that degraded model performance. We therefore applied a cross-dataset filter, retaining only the 2,360 genes detected as expressed in two independent K562 CRISPRi screens12,14, and added GFI1B as a critical test case (2,361 total, $\sim$21 million parameters). The effect was immediate: with reduced noise, the learning curve improved steadily throughout training, reaching $r=0.64$ with training r=0.65r=0.65—indicating minimal overfitting.
The contrast is stark: nearly quadrupling the gene count degraded performance by 42%, while filtering to cross-dataset reproducible genes yielded a 73% improvement. Notably, the larger model ($\sim$54 million parameters) underperformed the smaller one ($\sim$21 million parameters), the opposite of what would be expected if model capacity were the limiting factor. This establishes a counterintuitive but practically important principle: gene set quality, not quantity, determines the resolution of the AI microscope. Cross-dataset reproducibility filtering removes genes whose perturbation responses are dominated by technical or biological noise, enabling the model to learn sharper regulatory relationships.
@@
To evaluate how well CDT-II captures the trans-regulatory “fingerprint” of each perturbation, we averaged predictions and observations across cells for each held-out gene and computed Pearson $r$ on these mean trans-effect profiles: GFI1B ($r=0.88$), TNFSF9 ($r=0.86$), TFRC ($r=0.85$), CD44 ($r=0.85$), and CD52 ($r=0.75$), yielding a mean $r$ of 0.84. This high pseudo-bulk correlation demonstrates that CDT-II faithfully captures the transcriptome-wide pattern of perturbation effects—the regulatory “fingerprint” that distinguishes one perturbation from another.
To contextualize these correlations, we examined the empirical reproducibility of CRISPRi perturbation effects across independent studies. Morris et al. found that only 56% of GFI1B trans-targets identified in their STING-seq screen replicated in the independent Gasperini dataset12,14, and direct comparison of pseudo-bulk effect sizes between the two K562 CRISPRi studies yields $r\approx 0.79$. CDT-II’s pseudo-bulk correlation (mean $r=0.84$) surpasses this empirical ceiling, indicating that the model captures essentially all reproducible biological signal. The lower cell-level correlation ($r=0.64$) reflects stochastic variation inherent to single-cell measurements—noise that is averaged out at the pseudo-bulk level but represents a fundamental limit of single-cell prediction.
Trans-target prediction provided a stringent test of CDT-II’s ability to capture regulatory wiring. Among the top 50 predicted trans-targets, 26 overlapped with the top 50 experimentally determined GFI1B targets (52%, approximately 25-fold enrichment over the 2.1% expected by chance), and CD52 showed comparable overlap (25 of 50, 50%). Notable GFI1B trans-targets among these include G0S2, TXNIP, and JUN—genes involved in cell cycle arrest, oxidative stress response, and transcriptional regulation, respectively, consistent with GFI1B’s known functions in hematopoietic differentiation.
@@
Figure 3: CDT-II predicts cell-level perturbation effects. (A) Overall scatter plot of predicted versus actual log2 fold changes across all validation cells (n=2,037n=2{,}037 cells $\times$ 2,361 genes), showing Pearson $r=0.64$. Each point represents a cell-gene pair, with density indicated by color intensity. (B) Per-gene prediction performance for five validation genes (GFI1B, TNFSF9, TFRC, CD44, CD52), with pseudo-bulk correlations (mean across cells) ranging from 0.75 to 0.88 (mean $r=0.84$). (C) Trans-effect profile for GFI1B: mean predicted versus mean actual effects across all 2,361 genes, showing strong recovery of the perturbation-specific regulatory signature ($r=0.88$).
@@
The attention analysis follows the central dogma flow. DNA self-attention (Fig. 4A) reveals how the model learns relationships between genomic positions within the 115 kb window centered on the perturbation site. DNA-to-RNA cross-attention (Fig. 4B) shows how each gene attends to specific genomic regions, capturing transcriptional regulatory relationships. RNA self-attention (Fig. 4C–D) captures co-regulatory relationships between genes, revealing the gene regulatory network structure.
The RNA self-attention matrix $2{,}361\times 2{,}361$] captures learned co-regulatory relationships between genes. Extracting the GFI1B query row—representing which genes GFI1B “attends to” during inference—revealed a biologically coherent ranking. Cell cycle regulators including CDCA8, CDC20, KIF2C, and KIF14 appeared in the top ranks, consistent with GFI1B’s established role in cell cycle control during hematopoietic differentiation15.
Visualizing the top 5% of attention edges as a network (Fig. 5A) reveals GFI1B as a hub gene with multiple connection types: downstream targets (genes that GFI1B regulates), upstream regulators (genes that influence GFI1B), and bidirectional relationships suggesting reciprocal regulatory dynamics. This bidirectional structure captures the feedback loops characteristic of transcription factor networks.
To quantify the correspondence between attention-derived and experimentally determined regulatory networks, we ranked all 2,361 genes by GFI1B attention weight and, independently, by absolute experimental effect size (mean $|\log_{2}\text{FC}|$ upon GFI1B perturbation12), then measured overlap between the top $N$ genes from each ranking. At top 100, 28 genes appeared in both lists, representing a 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$ by hypergeometric test). This demonstrates that CDT-II’s attention maps recover known regulatory relationships without supervision.
@@
Figure 4: Attention maps reveal regulatory structure. (A) DNA self-attention patterns across the 115 kb genomic window for two layers. (B) DNA-to-RNA cross-attention profiles for five TSS perturbations.
@@
Figure 4 (continued). (C) RNA self-attention per-head analysis for GFI1B, showing that different attention heads capture distinct aspects of the regulatory network. (D) RNA self-attention 2D heatmap showing gene-gene co-regulatory relationships centered on GFI1B and its top-attended genes. (E) Cell-to-cell attention variation. Left: pairwise correlations of GFI1B attention vectors for 20 representative cells sampled from 477 validation cells (mean $r=0.75$ across all pairs), showing that core network structure is preserved while attention weights are modulated by each cell’s expression state. Right: genes with the most variable attention across cells, representing cell-state-dependent regulatory targets.
@@
We applied Louvain community detection16 to networks derived from each mechanism separately. From the RNA self-attention matrix, we identified community C1 (679 genes) with highly significant enrichment for RNA binding (Gene Ontology17 molecular function; $P=5\times 10^{-16}$) and metabolism of RNA (Reactome18; $P=2\times 10^{-9}$). From the DNA-to-RNA cross-attention matrix—where gene similarity was computed based on shared attention to the same genomic regions—we identified community C3 (165 genes) with even stronger enrichment for RNA binding ($P=1\times 10^{-16}$), spliceosome (KEGG19; $P=4\times 10^{-10}$), and mRNA splicing (Reactome18; $P=2\times 10^{-9}$).
The convergence between these independently derived modules was striking: 132 of 165 cross-attention genes (80%) were contained within the 679-gene self-attention community (2.8-fold enrichment over the 47.5 expected by chance; hypergeometric $P=9.3\times 10^{-46}$). This core module of 132 genes spans multiple levels of RNA processing: spliceosome components (SF3A3, SF3B4, PRPF3, SNRNP27, SNRPG, BCAS2, RBM8A), RNA modification and decay factors (YTHDF2 for m6A reading, ADAR for RNA editing, EXOSC10 for exosome-mediated degradation, TUT4 for uridylation, XPO1 for nuclear export), chromatin regulators (HDAC1, KDM1A/LSD1, ARID1A of the SWI/SNF complex), transcription machinery (CDC73 of the PAF complex, TAF12 of TFIID, TRIM33), and translation initiation factors (EIF2B3, EIF2B4, EIF4G3).
@@
That both mechanisms independently identify the same RNA processing module indicates that these genes are coordinately regulated at two distinct levels of the central dogma: they share transcriptional regulatory inputs (detected by cross-attention) and participate in the same post-transcriptional feedback networks (detected by self-attention) (Fig. 5B–C). Notably, while the architecture provides separate attention mechanisms, this convergence was not explicitly encouraged—it emerged from training on perturbation data alone.
@@
C
Figure 5: GFI1B regulatory network and convergent attention on an RNA processing module. (A) Network visualization of GFI1B regulatory relationships derived from the top 5% of RNA self-attention edges. GFI1B (red, center) functions as a hub gene connected to downstream targets (blue), upstream regulators (orange), and genes with bidirectional relationships (purple). Edge thickness is proportional to attention weight. (B) GO enrichment comparison showing both attention mechanisms identify “RNA Binding” as the top enriched term. RNA self-attention community C1 (679 genes) and DNA-to-RNA cross-attention community C3 (165 genes) show highly significant enrichment for RNA processing functions. (C) Venn diagram showing 132-gene core overlap (80% of cross-attention community). Core genes span spliceosome components, RNA processing factors, chromatin regulators, and transcription machinery.
Cross-attention recovers known regulatory elements
To test whether CDT-II’s cross-attention learns genuine regulatory structure, we compared attention patterns against ENCODE regulatory annotations20 for K562 cells (Fig. 6).
Enformer3, which provides CDT-II’s DNA embeddings, was pre-trained on epigenomic tracks including DNase-seq and histone ChIP-seq, so the initial embeddings already encode regulatory element positions. However, cross-attention does not operate on these raw embeddings: a projection layer and two DNA self-attention layers, all trained end-to-end on perturbation prediction, substantially transform the representations before cross-attention computes gene-specific regulatory maps (see Discussion).
For each of the five held-out genes, we extracted the mean cross-attention profile across validation cells, identifying the top 10% of attended genomic bins (89 of 896 bins spanning the 115 kb window). We then overlapped these high-attention bins with ENCODE K562 peak calls for five regulatory marks: DNase I hypersensitivity (open chromatin), CTCF binding, H3K27ac (active enhancers), H3K4me1 (poised enhancers), and H3K4me3 (active promoters). Fisher’s exact tests with Haldane correction were applied to each gene–mark combination (24 testable of 25; one combination, CD44 $\times$ H3K27ac, yielded no ENCODE peaks in the window).
[…]
To assess statistical robustness, we performed circular permutation tests (n=1,000n=1{,}000), shifting attention profiles along the genomic axis while preserving their autocorrelation structure. Among the 15 combinations tested (DNase, CTCF, and H3K4me3 $\times$ 5 genes), 14 achieved permutation $P<0.005$ (Supplementary Fig. 1a), confirming that the observed enrichments are not artifacts of attention smoothness or genomic feature clustering. Threshold robustness analysis showed that enrichments remained significant across a range of attention thresholds (top 5–20%), with effect sizes varying smoothly rather than appearing only at a single arbitrary cutoff (Supplementary Fig. 1b).
To further characterize what cross-attention captures, we classified each genomic bin by its ENCODE annotation—promoter (H3K4me3+), active enhancer (H3K27ac+ & H3K4me3-), CTCF-only, or unannotated—and compared attention weights across classes (Fig. 6B). Bin class separation was substantial: CTCF bins received significantly higher attention than unannotated bins (Cohen’s $d=2.4$), as did promoter bins ($d=1.4$). Kruskal–Wallis tests confirmed highly significant differences across all bin classes ($P=1\times 10^{-45}$). Notably, CTCF-associated bins showed the most consistent gene-specificity signal, ranking as the most informative mark across all five genes by entropy analysis.
These results demonstrate that CDT-II’s task-driven training selects and organizes genomic information into gene-specific regulatory maps. The cross-attention mechanism, optimized solely to predict perturbation effects, converges on the same regulatory grammar catalogued by ENCODE—providing strong evidence that the AI microscope reveals genuine regulatory biology rather than statistical artifacts.
Figure 6: Cross-attention recovers known regulatory elements. (A) Enrichment heatmap showing odds ratios (Fisher’s exact test) for overlap between CDT-II’s top 10% cross-attention bins and ENCODE K562 regulatory peaks across five held-out genes and five marks. 23 of 24 testable combinations achieve $P<0.001$ (CD44 $\times$ H3K27ac yielded no peaks). (B) Violin plots of attention weights stratified by ENCODE bin class (promoter, active enhancer, CTCF, unannotated); black lines indicate means, red lines indicate medians. CTCF-associated bins show the largest separation from unannotated bins (Cohen’s $d=2.4$), with Kruskal–Wallis $P=1\times 10^{-45}$ across all classes. See Supplementary Fig. 1 for circular permutation tests and threshold robustness analysis.